OpenStudy (anonymous):

A rectangular prism is truncated removing 20% of its height. The original height of the prism was 100 inches and it has a length of 5 inches and a width of 10 inches. What is the volume of the truncated prism?

OpenStudy (anonymous):

2,000 in^3 3,000 in^3 4,000 in^3 5,000 in^3

OpenStudy (cwrw238):

ht of truncated prism = 80% of 100 = 80 ins vol = w * h * l = 5 * 10 * 80 = 4000 cu ins